OpenStudy (anonymous):

write (square root) of -32 in simplified complex form

zepdrix (zepdrix):

Hey there :) Here is a clever way that we can break up the negative sign and the 32.\[\large\rm \sqrt{-32}\quad=\sqrt{-1\cdot 32}\quad=\sqrt{-1}\cdot\sqrt{32}\]

zepdrix (zepdrix):

Recall that we define our `imaginary unit` in this way: \(\large\rm \color{orangered}{\sqrt{-1}=i}\)

zepdrix (zepdrix):

Do you understand how that is going to help us?\[\large\rm \color{orangered}{\sqrt{-1}}\cdot\sqrt{32}\quad=\quad ?\]

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\large\rm \color{orangered}{\sqrt{-1}}\cdot\sqrt{32}\quad=\color{orangered}{i}\cdot\sqrt{32}\]Good, we can input our imaginary unit, ya? Another good step we can apply, is to pull anything out the root if we can.

zepdrix (zepdrix):

So does the value 32 contain any `factors` that are perfect squares? These are perfect squares: 4, 9, 16, 25, 36, ... Does 32 have any of those values as factors?

OpenStudy (anonymous):

ya 16 and 2

zepdrix (zepdrix):

Ok great, let's apply that earlier trick again,\[\large\rm i\cdot \sqrt{32}\quad=i\cdot \sqrt{16\cdot2}\quad=i\cdot \sqrt{16}\cdot\sqrt{2}\]

zepdrix (zepdrix):

And then what's the last simplification step you can do? :)

zepdrix (zepdrix):

See it?

OpenStudy (anonymous):

square root 16

OpenStudy (anonymous):

i